The x component of positive direction of x? A is -26.2 m and the y component is +52.2 m. (a) What is the magnitude of A? (b) What is the angle between the direction of A and the (a) Number 58.4 (b) Number 35
. (a) How far and (b) at what angle (measured counterclockwise from east) would a bird fly in a straight line from the same starting point to the same final point? (a) Number (b) Number Units
Answer #1

1)

Direction, theta = arctan (52.2/-26.2) = - 63.35 degree

Direction, theta = 360 - 63.35 = 296.65 degree

2)

net horizontal distance

Dx = - 3.1 km

net vertical distance

Dy = 3.3 - 5.1 = - 1.8 km

a)

net displacement, D ^2 = Dx^2 + Dy^2

D^2 = 3.1^2 + 1.8^2

D = 3.585 km

b)

Direction  

x = arctan ( 1.8/3.1)

x = 30.14 ccw from west.

=======

do comment in case any doubt will reply for sure.. Goodluck